OpenStudy (anonymous):

x^2-2/3=0 How do I solve for x?

OpenStudy (anonymous):

Difference of squares:\[(x+y)(x-y)=x^2-y^2\]So, factoring x^2 - 2/3 = 0 through difference of squares gives us:\[(x+\sqrt {2/3})(x-\sqrt {2/3}) = x^2-2/3=0\]\[x=\pm \sqrt{2/3}\]

OpenStudy (anonymous):

@revernt

OpenStudy (anonymous):

multiply by 3 and group like term ie 3x^2=2 and result is +or- square root of 2/3

OpenStudy (anonymous):

thanks!